Given a cost of $180,000 and a salvage value of $30,000, using the Straight-Line method what would the Book Value be at the end of the 7th year given a useful life of 15 years?

Answer by rfer(16322)

180000-30000=150000/15=10000 per yr
(150000-70000)+30000=$110,000